

********************************************************************
UARS WEEKLY STATUS REPORT - 19 FEB 1999
********************************************************************
***********
SOLSTICE
***********
SOLSTICE continues to operate normally, collecting solar data on all
available orbital tracking periods and stellar calibration observations,
whenever possible.
***********
SUSIM
***********
SUSIM successfully gathered its normal solar, offset, and occultation
data after Feb. 12 when it was powered on again without apparent
problems.
Reprocessing of the SUSIM data to data version 20 continues.
After several weeks of work and the generation of a new set of level 0
test files, SUSIM production software has passed all of the tests
for Y2K compliance. Our thanks go to Ed Esfandiari of the SUSIM group
as well as John Rizzello and N. Downs of the CDHF for their efforts.
***********
PEM
***********
PEM is active PMON control. Energy processor #6 for the HEPS-1 unit
has been found to not activate correctly. Data affected are from the upper
energy EE channel for Head-2 of HEPS-1 (that is the +15 deg electrons above
300 keV). Tests are being conducted on this energy processor to assure
that it is operating correctly. Measurements from this sensor will be
activated as soon as possible. The PEM team and FOT are currently examing
the PMON activation sequence to determine if there is interference during
the commanding of this energy processor.
While under PMON control, the PEM instrument computes energy deposition
from the HEPS unit.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
The instrument is now off as scheduled. MLS is scheduled to begin south viewing
operations on Feb 27 and operate though Mar 4.
***********
HRDI
***********
Nothing received.
***********
WINDII
***********
WINDII is currently off, and scheduled for operation March
5 to March 16. Work on the activation of the new production
software is under way.
***********
MPG
***********
Routine processing of instrument Daily Activity Plans continues
with no problems. The MPG continues to assist the Project
Scientist with the scheduling of instrument operations.
***********
CDHF
***********
The CDHF processed current day data through level 3 for all active
instruments. Additionally, CDHF personnel reprocessed: (R)= Requested
(C) = Complete, (F)= Failed (L) = Days remaining to reprocess.
0 days HAL (4.80) PROD (L3) (1874R-1721C-1F=152L)
34 days PEM (3.32) PROD (OA-L2) (760R-84C=626L)
2 days WIN (5.11) ENG (L1-L3) (2R-2C=0L)
1 day SOL (4.62) PROD (1R-1C=0L)
The CDHF will load December 1998 current day DAAC data in February 1999.
CDHF pending software upgrades: DEC AXP Fortran 6.5, DEC AXP OPEN
VMS 6.2, OpenIngres.
***********
FOT
***********
This report covers 05 February 1999 (Orbit 40467, GMT Day 037) through 12
February 1999.
SPACECRAFT OPERATION
The observatory is now in Reverse Flight (Northern Hemisphere viewing) and
is performing nominally under two-battery operations. The instrument
operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
ACRIM 038/2344 40496 Off
PEM 038/2344 40496 Off
SUSIM 038/2344 40496 Off
WINDII 038/2344 40496 Off
SOLSTICE 039/0130 40497 On-PMON
HALOE 039/1142 40503 On
MLS 042/2244 40555 Off
ACRIM 042/2357 40556 On
SOLSTICE 043/0007 40556 On-Full
HRDI 043/0007 40556 On-Full
SUSIM 043/1339 40564 On
The beta angle ranged between the angles of 65.1 degrees to 34.1 degrees
for this report period. The beta angle is now decreasing toward a minimum
angle of -2 degrees for the next yaw on 20 February 1999 (DOY 051).
Spacecraft battery 1, 2, and 3 performance monitoring continues. Battery 1
remains OFF the charge relay. Battery 2 and Battery 3 are currently
maintaining greater than 23.8 V end-of-night (EON) load bus voltage.
Battery temperatures are stable with a temperature delta between Battery 2
and Battery 3 of 0.9 to 1.1 degrees C. Additional battery performance data
for each battery on 05 February 1999 and 12 February 1999 are provided in
Appendix D.
The clock error ranged from -13.5 to 11.4 msec during this report period.
Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT for
this report. A summary is listed below and the details of ALERT and Out-
Of-Limits occurrences are listed in Appendix B.
% ALERTS = 00
% OUT-OF-LIMITS = 17
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
% FOT 04
% ACRIM 02
% SUSIM 02
% HALOE 01
% MLS 01
% WINDII 01
% PEM 01
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
% FOT 01
One (1) Anomaly Report remained OPEN at the end of this report period (see
Appendix E).
TRANSPONDER FREQUENCIES:
Transponder A: 99/019 - 2287.496951
Transponder B: 99/020 - 2287.499045
GROUND SYSTEM OPERATION
One (1) operational support problems occurred this period. Anomaly Reports
are included in Appendix E.
The data loss calculations supplied by Data Capture Facility (DCF) from
mission start through UTC Day 99/035 show the observatory data loss to be
21 hours, 50 minutes, 51.540 seconds. This is a 0.0338 percent data loss
which equals a 99.9662 percent data capture for the mission.
There were no data loss calculations reported for this report period due to
a power outage in Building 23. The outage lasted from approximately
042/1200 to 043/2240.
The FOT received a partial set of test 8-week planning aids last week. The
files start in 1999 and cross into year 2000. These planning aids were
processed by the FOT this week. The FOT are still waiting for the
remainder of these test products from FDF.
The UARS Test and Training Simulator (UTTS) remains non-operational. The
estimated time to repair the UTTS and return it to operation is unknown.
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
The FOT continues to support the PEM team with diagnosis of the Energy
Processor 6 status error.
The FOT continues to investigate the feasibility of a new PMON action to
change battery V/T curves based on loading to mitigate battery risks in sun
pointing safehold.
YEAR 2000 OPERABILITY SUPPORT - No change
UARS PROPULSION MODULE B/D THRUSTER LINE FILL
Planning continues for the simulator testing.
UARS RE-ENGINEERING - TELEMETRY PROCESSING - No change
MEETINGS
None.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw Around Maneuver is scheduled for 20 February 1999.
The next annular solar eclipse will occur on 16 February 1999 over the
Indian Ocean, Australia and the South Pacific Ocean.
PERSONNEL CHANGES
None
Attachments: Appendix A, Real-time Commands Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 12 February 1999
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ -------
40476 037/1512 99-054 OBC Clk Adj fr 41.2 to 20.6 msec/day
40483 038/0303 99-055 OBC Clk Adj fr 20.6 to 0.00 msec/day
40496 038/2344 99-056 ACRIM Power OFF
40496 038/2344 99-057 PEM AXIS OFF
40496 038/2344 99-058 SUSIM Power OFF
40496 038/2344 99-059 WINDII Power OFF
40503 039/1142 99-060 HALOE Power ON
40505 039/1416 99-061 OBC Clk Adj fr 0.00 to 20.6 msec/day
40541 041/2355 99-062 OBC Clk Adj fr 20.6 to 0.00 msec/day
40555 042/2244 99-063 MLS Power OFF
40556 042/2357 99-064 ACRIM Power ON
40564 043/1339 99-065 SUSIM Power ON
_______________________________________________________________
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 12 February 1999
MNEMONIC STATE DESCRIPTION ORBITS
--------- ----- ------------ ------
WIQBCURRENT Y-LO WINDII Quiet Bus Current Continuous
SOSDTCTTMPR Y-LO Detectors Temperature 40509-15,22-25
SORLGCTMPR Y-LO Logic Temperature 40509,10,12-14,22-25
SOSTMPRBLFT Y-LO Temperature Bot_Left_Front 40514,23
QTSSPPMX Y-HI SSPP -X Thermostat 40520,23,33,34,48,52,53
ACTEMPANA1 Y-LO Temp PCU Board 40556-62
ACTEMPANA2 Y-LO Temp Sensor Electronics 40556-62
ACTEMPSA1 Y-LO Temp ACR 1 40556-62
ACTEMPSA2 Y-LO Temp ACR 2 40556-62
ACTEMPSA3 Y-LO Temp ACR 3 40556-62
SORLBXTMPR Y-LO Logic_Box Temperature 40556-62
SOSTMPRTRFT Y-LO Temperature Top_Rt_Front 40556-62
SOSTMPRTRBK Y-LO Temperature Top_Rt_Back 40556-62
SOSTMPRTLMD Y-LO Temperature Top_Left_Mid 40556-62
SOSTMPRTLBK Y-LO Temperature Top_Left_Back 40556-62
SOSDTCTTMPR Y-LO Detectors Temperature 40556-62
SOSTMPRBLMD Y-LO Temperature Bot_Left_Mid 40556-62
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
--------- ------------ ------
None this report period
_______________________________________________________________
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 12 February 1999
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ---- -----------------
40476 037/15:15 (06 Feb 99) from 41.2 to 20.6 msec/day
40483 038/03:08 (06 Feb 99) from 20.6 to 0.00 msec/day
40505 039/14:20 (08 Feb 99) from 0.00 to 20.6 msec/day
40541 041/23:58 (10 Feb 99) from 20.6 to 0.00 msec/day
_______________________________________________________________
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 12 February 1999
05 Feb 1999 - GMT Day 036 Beta = 65.1 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ------- ------- ------
1 +11.2/-5.6 -3.68/-4.94 +0.0/ +0.0 16.8 0.000 0.0
2 +22.4/-11.2 +3.87/+2.92 +22.8/-16.0 26.4 1.027 18.9
3 +56.0/-11.2 +2.92/+1.97 +22.4/-15.2 26.2 1.030 18.5
12 Feb 1999 - GMT Day 043 Beta = 34.1 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ------- ------- ------
1 0.0/-11.2 -3.05/-4.31 +0.0/ +0.0 16.8 0.000 0.0
2 +56.0/ 0.0 +5.47/+4.51 +30.4/-14.4 26.2 1.030 21.1
3 +56.0/-11.2 +4.51/+3.55 +29.6/-14.4 26.1 1.027 20.8
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
----- -------- ----------
no changes V/T 5
APPENDIX E
UARS ANOMALY REPORTS
Week ending 12 February 1999
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
99-006 FOT 40469 037/0439 14170 Generic Late Acq
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
None this report period
ANOMALY INVESTIGATION REPORTS GENERATED
AIR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------- ------ ----- ---- --- -----------------
None this report period
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
99-006 FOT see list Feb 99 14170 Generic Late Acq
SEVNTFW99/07